OpenStudy (anonymous):

limit as n approach infinity for the summation of 1/[(n+3)(n+4)] from n=1 to infinity

OpenStudy (turingtest):

I think you have written this funny

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \sum_{n=1}^{ \infty}1/[(n+3)(n+4)]\]

OpenStudy (anonymous):

is it \[\sum_{n=1}^{\infty}\frac{1}{(n+3)(n+4)}\]?

OpenStudy (anonymous):

ok gimmick is partial fractions rewrite at \[\frac{1}{n+3}-\frac{1}{n+4}\] and you will see that your sum "telescopes"

OpenStudy (anonymous):

i want to know whats next after we partialise

OpenStudy (turingtest):

observe how the terms behave, try writing out the first 3 or so...

OpenStudy (anonymous):

ok sure, but first of all the thing you wrote above doesn't really make any sense it is \[\sum_{n=1}^{\infty}a_n:=\lim_{k\to \infty}\sum_{n=1}^ka_n\]

OpenStudy (anonymous):

do what mr @TuringTest said. replace n by 1, by 2, by 3 etc and you will see exactly what happens since the second term is minus the previous one, you will have a raft of cancellation, leaving only a term or two check it

OpenStudy (anonymous):

do you mean to say that the terms b4 partialising are different from those aftr partialising

OpenStudy (turingtest):

what are the first three terms in the series after applying partial fractions?

OpenStudy (anonymous):

1/20+1/30+1/42.......

OpenStudy (turingtest):

no, do it in the way that satellite showed you in terms of the representation in partial fractions

OpenStudy (turingtest):

\[\frac1{n+3}-\frac1{n+4}\]what are the first three terms?

OpenStudy (turingtest):

...or however many you feel like writing

OpenStudy (anonymous):

[1/4-1/5]+[1/5-1/6]+[1/6-1/7]+[1/7-1/8].......

OpenStudy (turingtest):

nice, now are you seeing some things cancelling?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

what is left?

OpenStudy (anonymous):

the first and the last term

OpenStudy (turingtest):

but what is the last term as n goes to infinity?

OpenStudy (anonymous):

it is infinity

OpenStudy (turingtest):

nice try, but infinity is not a number, so the answer is that there \(is~no~last~term\) for whatever term we call the last we can always make another that will cancel the one before it, hence they will \(all\) cancel except the first term the logic here may be a bit counter-intuitive at first, but remember 1) infinity is not a number; it is a concept 2) there is no last term, we can always add another to cancel the one before it

OpenStudy (turingtest):

...meditate on this, it will help you develop a mathematical intuition ;)

OpenStudy (anonymous):

okay, so we are only left with the first term ?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

\[(\frac14-\cancel{\frac15})+(\cancel{\frac15}-\cancel{\frac16})+(\cancel{\frac16}-\cancel{\frac17})+...\]you can see the pattern of cancellation continues literally forever, so no last term remains, only the first

OpenStudy (anonymous):

okay i get that

OpenStudy (turingtest):

okay, so what does that tell you about\[\sum_{n=1}^\infty\frac1{(n+3)(n+4)}=\sum_{n=1}^\infty\frac1{n+3}-\frac1{n+4}\]?

OpenStudy (anonymous):

that the limit approach approach the first term

OpenStudy (turingtest):

no limit, the sum of the series is 1/4 (it's not the first term, which is 1/4-1/5)

OpenStudy (turingtest):

now here is where your question makes no sense\[\sum_{n=1}^\infty\frac1{n+3}-\frac1{n+4}=\frac14\]but you typed the problem as\[\lim_{n\to\infty}\sum_{n=1}^\infty\frac1{n+3}-\frac1{n+4}=\lim_{n\to\infty}\frac14=\frac14\]so the limit thing is redundant and adds nothing to the problem the sum of the series is a constant, and the limit of a constant is a constant, so why the hell is that even in there if you made a typo that would make sense, for instance if it were....

OpenStudy (turingtest):

\[\lim_{n\to\infty}\sum_{i=1}^n\frac1{i+3}-\frac1{i+4}=\lim_{n\to\infty}\frac14=\sum_{i=1}^\infty\frac1{i+3}-\frac1{i+4}=\frac14\]that would not be redundant as it would be talking about the limit of the highest term in the series n approaching infinity same answer either way though, so who cares :P

OpenStudy (turingtest):

typo that was meant to read\[\lim_{n\to\infty}\sum_{i=1}^n\frac1{i+3}-\frac1{i+4}=\sum_{i=1}^\infty\frac1{i+3}-\frac1{i+4}=\frac14\]

OpenStudy (anonymous):

thnx